OpenStudy (anonymous):

Evaluate: sin2 25o + sin2 65o.

OpenStudy (anonymous):

its sin^2 25degree+sin^2 65degree

OpenStudy (midhun.madhu1987):

sin2 25o + sin2 65o = sin 65o can be written as cos(90 - 25)o and evaluate it !!!

ganeshie8 (ganeshie8):

trig identity : \[\large \sin (\theta) = \cos (90 - \theta)\]

OpenStudy (anonymous):

cos 90=0 and cos 25=?

ganeshie8 (ganeshie8):

write \(\large \sin(25)\) as \(\large \cos(90 - 25)\)

ganeshie8 (ganeshie8):

which is same as \(\large \cos(65)\)

OpenStudy (anonymous):

cos(90-65),cos(25)

OpenStudy (anonymous):

but sin^2

ganeshie8 (ganeshie8):

see if below makes sense : \[\large \sin^2 25 +\sin^2 65\] \[\large \left(\sin 25\right)^2 +\sin^2 65\] \[\large \left(\cos(90- 25)\right)^2 +\sin^2 65\] \[\large \left(\cos(65)\right)^2 +\sin^2 65\] \[\large \cos^2 25 +\sin^2 65\] \[\large 1 \]

OpenStudy (anonymous):

65+65=1 25+65=1?

ganeshie8 (ganeshie8):

typo in second line from last : \[\large \large \cos^2 \color{Red}{65} +\sin^2 65\]

ganeshie8 (ganeshie8):

another trig identity : \[\large \cos^2\theta + \sin^2\theta = 1\]